OpenStudy (anonymous):

Solve for x 2^x+2^4-x=17

OpenStudy (anonymous):

Is is 2^(4-x)?

OpenStudy (anonymous):

0 & 4

OpenStudy (anonymous):

@ Abe le Babe: No

OpenStudy (anonymous):

no!!is it (2^4)-x?

OpenStudy (anonymous):

2"x" + 2"4-x" = 17

OpenStudy (anonymous):

Does that make more sense?

OpenStudy (anonymous):

That's better...:-)

OpenStudy (anonymous):

@ Sritama: How do you get to that

OpenStudy (anonymous):

Probably by eye...

OpenStudy (anonymous):

\[2^x + 2^{4-x} = 2^4 + 2^0\] Does that help?